A reaction has ΔH∘rxn=−237 kJ and ΔS∘rxn= 218 J/K. Calculate ΔG∘rxn at 60 ∘C.
Answer #1

Assume there is no chang ein H and S with respect of T

T = 60°C = 60+273 = 333 K

then

G = H - TS is valid

H = -237000 J

S = 218 J/K

then

G = -237000 -333*218 = -309594 J

G = -309.59 kJ